You hang two pendulums next to one another. They both are 3 meters long.

One of them has a 1 kg mass at the bottom, the other has a 2 kg mass at

the bottom. You pull the lighter ball back so that it forms a 40◦ angle with

the vertical, and then let it go. It collides elastically with the other pendulum.

What is the maximum angle the heavier pendulum creates with the vertical?
Answer #1

Line AB is the minimum gravitational potential energy line on which we define the potential of any mass to be zero .

Along the line AB :

Just Before collision : Only 1kg ball is moving with velocity v1

Just after collision : 3 kg ball is moving with velocity v2 and 1 kg with v1 ( but in opposite direction )

-444444 360540° 3 coso 3m to 3m 12 from the figure above, we can see 3= h + 3 Los 400 h = 3 - 3 cos 40° = 3-2.298 h = 0.702 mLet, just after collision, the velocity of 3 kg mass be v₂ So, Conservating linear momentum along the line AB Initial momentucoefficien for the of where, e case of Elastic collision, Restitution , e=1 e is defined as velocity of seperation velocity oMaximum height ₂ 3 kg ball, when collision has been is achieved by the all its [(KE) just after converted into Potential Ener